OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (tkhunny):

What is Slope-Intercept Form? Where do we get the Slope information? Why did they give us some other line? What's the big idea giving us a point that is not the y-intercept?

OpenStudy (anonymous):

y=mx+b

OpenStudy (tkhunny):

Okay, how about the two additional questions?

OpenStudy (anonymous):

y=-1/4x=3 by my calculations. Hope this helps! Lizzy